OpenStudy (anonymous):

can someone solve this please! using L'Hospitals rule (x-ln(x)) as x approaches infinity thank you

OpenStudy (anonymous):

x grow a lot faster than ln(x) so the limit is +Infinity

OpenStudy (cruffo):

Can't apply L'H directly since the difference in an indeterminate form. There is What if you study the ratio of \[\frac{x}{\ln x}\] If this approaches infinity as x approaches infinity then you know that x grows faster, so the original problem has a limit of infinity as well.

OpenStudy (anonymous):

You can apply L'Hospital's rule by noticing \[ \frac{(x-\ln (x)) (x+\ln (x))}{x+\ln (x)}=\frac{x^2-\ln ^2(x)}{x+\ln (x)} \]

OpenStudy (anonymous):

by taking the derivative of up and down, you get \[ \frac{2 x - 2 \frac{\ln (x)}{x}} {1+ \frac 1 x} \] You can see that the above ratio goes to \(+\infty\)

OpenStudy (anonymous):

ok so you multiplied the numerator and the denominator by (x+ln(x)) why did you do that?

OpenStudy (anonymous):

to be able to use L'Hospital's rule

OpenStudy (anonymous):

what is the requirements ba to use the L'Hospitals Rule @eliassaab

OpenStudy (anonymous):

@KingGeorge howdo you solve this

hartnn (hartnn):

the requirement to use L'Hospitals rule is that the numerator/denominator should be of the form 0/0 or infinity/infinity if u directly put the value of x

OpenStudy (anonymous):

ok then how about your final answer should it also be infinity or zero? @hartnn

hartnn (hartnn):

ok,the purpose of using L'Hospitals rule is to remove such indeterminant form. if removed by differentiation,then final answer would be finite

OpenStudy (anonymous):

oh so the answer of elissab earlier was correct? i just need to differentiate it more?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes,the answer is correct, u will get same answer if u diff more.......so no need

OpenStudy (anonymous):

so wait example that was given in an exam what should be my final answer? @hartnn

hartnn (hartnn):

infinity only.

OpenStudy (anonymous):

ohh ok thanks men so i will just sub infinity in the equation?

hartnn (hartnn):

if u diff it just once more(after simplifying),u will get constant in denominator.. then u substitute infi,to get final answer as infinity

OpenStudy (anonymous):

ohh ok thanks :)

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

can you solve me one last questioon before i leave?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

sure.

OpenStudy (anonymous):

ok i'll make another post wait ffor it

hartnn (hartnn):

ok